# Mutual Information Estimation via Score-to-Fisher Bridge for Nonlinear Gaussian Noise Channels

### v1

@@ Abstract
We present a numerical method to evaluate mutual information (MI) in nonlinear Gaussian noise channels by using denoisingscore matching learning for estimating the score function of channel output. Via de Bruijn’s identity and the I–MMSE relation, Fisher information estimated from the learned score yields accurate estimates of MI through an integral representation of MI for a variety of priors and channel nonlinearities (e.g., elementwise nonlinearity). In this work, we propose a comprehensive theoretical foundation for the Score-to-Fisher bridge methodology, along with practical guidelines for its implementation. We also conduct extensive validation experiments, comparing our approach with closed-form solutions and a kernel density estimation baseline. The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for mutual information estimation in nonlinear Gaussian noise channels.
@@ I Introduction
and study how to estimate the mutual information $I(X;Y_{t})$ efficiently and accurately. The nonlinear front-end can be represented by a deterministic function $f$. Such models arise widely due to nonlinear front-end saturation/clipping, DAC/ADC nonlinearity, photonic modulators, optical fibers, or learned encoders implemented with neural networks. The value of the mutual information $I(X;Y_{t})$ is important for evaluating the ultimate performance of coded communication systems and it provides useful information for designing the system such as coding, preprocessing, or decoding/detection algorithms.
When $f$ is nonlinear, evaluating $I(X;Y_{t})$ is, in general, challenging in practice because accurate posterior means $\mathbb{E}[X|Y_{t}]$ and normalized likelihoods are inaccessible or computationally intractable. For linear Gaussian channels, the de Bruijn identity and the I–MMSE relation yield
$\frac{d}{dt}I(X;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}=-\frac{1}{2t^{2}}\mathrm{mmse}(t),$ (2)
so that the mutual information $I(X;Y_{t})$ can be obtained by integrating either the Fisher information $J(Y_{t})$ or the MMSE. Our key observation is that the deterministic nonlinearity can be absorbed into a transformed input $W=f(X)$, so that
$I(X;Y_{t})=I(W;Y_{t}),\quad Y_{t}=W+Z_{t},$
holds for any deterministic function $f$. Hence, for general $f$ we can still rely on AWGN identities:
$\mathrm{mmse}_{W}(t)=nt-t^{2}J(Y_{t}),$
and consequently obtain $I(X;Y_{t})$ by integrating $J(Y_{t})$ or $\mathrm{mmse}_{W}(t)$.
In this paper, we propose a score-based estimator of mutual information tailored to $Y_{t}=f(X)+Z_{t}$. The method learns the score of channel output, or the marginal $p_{Y_{t}}$,
$s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}),$ (3)
using denoising score matching (DSM) and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an integral representation of MI:
$\widehat{I}(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-\widehat{J}(Y_{t})\Big)\,dt.$ (4)
Crucially, the approach never requires access to the posterior $p_{X|Y}(\bm{x}|\bm{y})$. All we need are forward samples $\bm{x}\sim p_{X}$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition, which enables a single-loop estimator of MMSE and the Fisher information.
@@
  • Score-only MI estimation for nonlinear front-ends. We formalize a pipeline— denoising score matching of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ MMSE/MI integration—that applies to arbitrary deterministic $f$ as soon as $p_{Y_{t}}$ can be sampled.
  • Unified theoretical grounding. Leveraging de Bruijn and I–MMSE, we derive $\mathrm{mmse}_{W}(t)=nt-t^{2}J(Y_{t})$ and connect $I(X;Y_{t})$ to $I(W;Y_{t})$ with $W=f(X)$, extending linear-Gaussian identities to general nonlinear front-ends without ambiguity.
  • Practical recipe. We detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral, and a diagnostic “double estimate” that compares a Tweedie-based direct MMSE with the Fisher-based MMSE.
  • Validation and use-cases. We verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
In linear Gaussian channels, I–MMSE and multivariate de Bruijn provide differential and integral representations of information quantities have been extensively studied. For nonlinear $f$, however, posterior-free and normalization-free MI estimators remain limited. Building on score matching and denoising-based score learning, our approach attaches directly to the marginal $p_{Y_{t}}$, circumventing posterior computations and partition functions. The estimate of mutual information can be used many applications such as design of coding systems, optimization of preprocessing, and design of detection and decoding algorithms.
@@ II-A Notation
$J(U)\equiv\mathbb{E}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right]\quad\text{with }\bm{u}\sim p_{U}$ (5)
@@ II-B Related works
For channels with memory, simulation-based methods have been developed to evaluate information rates without closed-form expressions. Arnold et al. [3] proposed a simulation-based framework that computes upper and lower bounds on the information rate by combining auxiliary finite-state channel models with forward–backward recursions and importance sampling; the method is broadly applicable to ISI and finite-state channels where exact evaluation is infeasible. Building on this line, Sadeghi et al. [4] optimized these bounds over the choice of auxiliary models, yielding systematically improved estimates and a practical tool for design and benchmarking of channels with memory.
Another line of work estimates $I(X;Y)$ directly from its definition via simulation-based plug-in estimators: one approximates $p_{Y}$ (or $h(Y)$) using Kernel Density Estimation (KDE) or mixtures [5] or employs $k$-nearest neighbors (kkNN) estimators [6]. In the additive Gaussian setting, recent theory shows that plug-in estimation of $h(X+Z)$ is in fact optimal [7].
@@ II-C AWGN channel and score function
where $t$ represents the noise variance. Let $p_{Y_{t}}$ denote the marginal density of $Y_{t}$. The score function of $p_{Y_{t}}$ is defined by $s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}).$ Throughout, we assume that sampling from $p_{X}$ is computationally feasible by forward sampling or Langevin sampling using the score $\nabla_{\bm{x}}\log p_{X}(\bm{x})$ (e.g., Rossky et al. [8], Welling and Teh [9]).
@@
$J(Y_{t})\equiv\mathbb{E}_{Y_{t}}\bigl[\|s_{Y_{t}}(\bm{y})\|^{2}\bigr].$ (7)
Under standard regularity conditions, the following Tweedie identity [10, 11, 12] holds for $t>0$:
@@ II-D I–MMSE relation
Let $h(Y_{t})$ denote the differential entropy of $Y_{t}$. Then the (multivariate) de Bruijn identity gives
@@
The de Bruijn identity was discussed in [15, 16].
@@ II-F Denoising score matching (DSM) method
where this equivalence was shown by Vincent [18]. This means that minimizing the DSM loss function is equivalent to minimizing the expectation of the squared distance between the learned score function and the true score function, which is sometimes called the Fisher divergence. This means that the unique minimizer of the DSM loss function satisfies $s_{\bm{\theta}}(\bm{y})\simeq\nabla_{\bm{y}}\log p_{\sigma}(\bm{y})$ in high probability. Equivalently, since $\bm{v}-\bm{w}=\sigma\bm{\epsilon}$, one may implement the loss function as $(\bm{v}-\bm{w})/\sigma^{2}$:
@@
In practice, we use a neural network to parameterize the score model $s_{\bm{\theta}}(\bm{v})$. Such a score model is also called a score network. A mini-batch training with stochastic gradient descent or other variants such as Adam is typically used to minimize the DSM loss function.
@@ III Estimating Mutual Information in AWGN Channels
In this section, we present a method to estimate the MMSE and MI for AWGN channels described by
@@ III-A MMSE–Fisher relation
A sketch of proof is the following. Using $I(X;Y_{t})=h(Y_{t})-h(Z_{t})$ with $h(Z_{t})=(n/2)\log(2\pi e\,t)$, we have
@@
by de Bruijn identity. Equating this with the I–MMSE form
$\frac{d}{dt}I(X;Y_{t})=-\frac{1}{2t^{2}}\mathrm{mmse}(t)$ (21)
gives $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$. ∎
Reeves et al. [21] investigated this identity in the context of linear Gaussian channels, highlighting its implications for non-Gaussian input distributions. Their work extends the I–MMSE framework beyond the classical Gaussian setting.
III-B Integral representations of mutual information
@@
$I(X;Y_{T})=$ $\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt$ (22)
$=$ $\frac{1}{2}\int_{T}^{\infty}\!\frac{\mathrm{mmse}(t)}{t^{2}}\,dt.$ (23)
@@
The integral representation directly follows by combining the de Bruijn identity [15] with the I–MMSE relation [13, 14].
@@ III-C Estimation of Fisher Information
Using DSM method, we can estimate the score function of $Y_{t}$. With $p=p_{X}$ and $\sigma^{2}=t$, the learned score model coincides with the score of $Y_{t}=X+Z_{t}$:
$s_{\bm{\theta}}(\cdot)\approx\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=s_{Y_{t}}(\bm{y}),$ (25)
where $s_{\bm{\theta}}(\cdot)$ can be implemented with a neural network with parameter vector $\bm{\theta}$. This enables estimation of $J(Y_{t})=\mathbb{E}\|s_{Y_{t}}(Y_{t})\|^{2}$ from samples of $\bm{x}\sim p_{X}$ and Gaussian noises.
@@
$=\mathbb{E}_{Y_{t}}\!\left[\left\|\,s_{\bm{\theta}}(Y_{t})-s_{Y_{t}}(Y_{t})\,\right\|^{2}\right]+\text{const},$ (27)
where the constant term does not depend on $\bm{\theta}$. Hence, with a sufficiently rich function class for $s_{\bm{\theta}}$ (e.g., a universal neural network) and sufficiently many samples, any sequence of empirical minimizers achieves $\mathbb{E}_{Y_{t}}\!\left[\left\|\,s_{\bm{\theta}}(Y_{t})-s_{Y_{t}}(Y_{t})\,\right\|^{2}\right]\rightarrow 0,$ i.e., the learned score converges to the true score in the Fisher divergence sense.
@@ III-D Gaussian input case for validation
In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$. It is well known [2] that $I(X;Y_{t})=(n/2)\log(1+P/t)$; see also [13] for the relation to MMSE.
@@
and $J(Y_{t})=\mathbb{E}\|s_{Y_{t}}(Y_{t})\|^{2}=n/(P+t)$. The remaining identities follow by substituting $J(Y_{t})$ into $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ and by $h(X\mid Y_{t})=h(Y_{t})-I(X;Y_{t})$ with $h(Y_{t})=(n/2)\log\big((2\pi e)(P+t)\big).$
@@ III-E Linear Gaussian channel case
IV Estimating MMSE and Mutual Information
Using the identity between $\mathrm{mmse}(t)$ and the Fisher information, we can evaluate $I(X;Y_{T})$ by numerical integration. In this section, we present details of the method to estimate the MMSE and MI for the AWGN channel family.
IV-A Score-based Estimation Method for MI and MMSE
@@
Algorithm 1 Score-to-Fisher bridge: Estimation Algorithm for MI and MMSE
1: Prepare a logarithmically spaced grid over $t$ (details of the log grid are described in Subsection IV-C).
@@
3: Denoising score matching (DSM) learning for $p_{Y_{t}}$: train the score model parameters by minimizing the empirical DSM loss:
@@
Any stochastic optimizer, such as Adam, can be applied to minimized the empirical loss function.
@@
$\widehat{J}_{t}\equiv\frac{1}{N}\sum_{i=1}^{N}\big\|s_{\theta}(\bm{y}_{i})\big\|^{2},$ (38)
@@
5: Estimating the MMSE:
$\widehat{\mathrm{mmse}}(t)=nt-t^{2}\widehat{J}_{t}.$ (39)
6: Estimating the MI via the integral representation:
@@
    where $\mathcal{T}$ is a training distribution over $t$ (e.g., logarithmically spaced as described later) and $\lambda(t)$ is a weight (e.g., $\lambda(t)=t$).
In the case of Scheme B, the score network has two inputs: the channel output $\bm{y}$ and the noise level $t$. In generative modeling frameworks for image generation, Scheme B is often used to train a noise-conditional score model. The training distribution $\mathcal{T}$ is logarithmically spaced as described in Subsection IV-C. The weight $\lambda(t)$ is a weight function that is used to weight the loss function.
IV-B Validating the Fisher-based estimate
Although Algorithm 1 presented in the previous subsection is an asymptotically consistent method to estimate the MMSE and MI, the accuracy of the estimate is not clear in the case of finite samples. One possible source of the error is the estimation error of the Fisher information due to inaccurate estimation of the score model. Although we can expect the accurate score matching in the Fisher divergence sense in the case of infinite samples, the convergence rate depends on the prior distribution of the source. We also have to consider the numerical error of the numerical integration.
In this section, we present a method to validate the estimated MMSE and MI. As a diagnostic purpose, we can compute the direct MMSE estimate based on the Tweedie formula:
$\widehat{\mathrm{mmse}}_{\rm dir}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\big\|\bm{x}_{i}-\{\bm{y}_{i}(t)+t\,s_{\theta}(\bm{y}_{i}(t))\}\big\|^{2},$ (42)
and check agreement with the Fisher-based estimate:
$\widehat{\mathrm{mmse}}_{F}(t)\equiv nt-t^{2}\widehat{J}_{t}.$ (43)
This consistency check is useful to validate the accuracy of the Fisher-based estimate.
IV-C Log-domain trapezoid integration
The integrand in the integration representation of MI is smooth and may have a long tail. The choice of the integration grid is critical for achieving high numerical accuracy. In this subsection, we present a geometric (log-spaced) grid for the integration. For $t\in[t_{\min},t_{\max}]$ with $M$ points, take $\log t$ to be equally spaced:

### v2

@@ Abstract
We present a numerical method to evaluate mutual information (MI) in nonlinear Gaussian noise channels by using denoising score matching (DSM) learning for estimating the score function of channel output. Via de Bruijn’s identity, Fisher information estimated from the learned score function yields accurate estimates of MI through a Fisher integral representation for a variety of priors and channel nonlinearities. In this work, we propose a comprehensive theoretical foundation for the Score-to-Fisher bridge methodology, along with practical guidelines for its implementation. We also conduct extensive validation experiments, comparing our approach with closed-form solutions and a kernel density estimation baseline. The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for MI estimation in nonlinear Gaussian noise channels.
@@ I Introduction
and study how to estimate the mutual information (MI) $I(X;Y_{t})$ efficiently and accurately. A nonlinear front-end function can be represented by a deterministic function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Such models arise widely due to nonlinear front-end saturation/clipping, DAC/ADC nonlinearity, photonic modulators, optical fibers, or learned encoders implemented with neural networks. The value of the mutual information $I(X;Y_{t})$ is important for evaluating the ultimate performance of coded communication systems; it provides useful information for designing the system such as coding, preprocessing, or decoding/detection algorithms. In a sensing system, the nonlinear front-end can be interpreted as a details of sensing process. The value of the mutual information $I(X;Y_{t})$ indicates how much information is transferred from the sensing target to the output.
When $f$ is nonlinear, evaluating $I(X;Y_{t})$ is, in general, challenging in practice because accurate posterior means $\mathbb{E}[X|Y_{t}]$ and normalized likelihoods are inaccessible or computationally intractable. Let $W=f(X)$ be a transformed input. The de Bruijn identity immediately yields
$\frac{d}{dt}I(W;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (2)
so that the mutual information $I(X;Y_{t})$ can be obtained by integrating the Fisher information $J(Y_{t})$.
Our key observation is that the Fisher information $J(Y_{t})$ can be estimated from the score function of $Y_{t}$ which can be estimated by denoising score matching (DSM) learning from samples of $Y_{t}$, and that the deterministic nonlinearity can be absorbed into a transformed input $W=f(X)$, so that
$I(X;Y_{t})=I(W;Y_{t}),\quad Y_{t}=W+Z_{t},$ (3)
holds for any deterministic function $f$. This equality is the basis of our score-based estimator of MI.
By integrating (2) over $t\in[T,\infty)$, for general $f$, we have the Fisher integral representation of MI:
$I(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt$ (4)
that depends only on the Fisher information of $Y_{t}$.
In this paper, we propose a score-based estimator of MI tailored to $Y_{t}=f(X)+Z_{t}$ [^1] A software implementation to reproduce the numerical results in this paper is available at https://github.com/wadayama/score-to-fisher-mi. The method learns the score of channel output, or the marginal density $p_{Y_{t}}$, using DSM learning and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an Fisher integral representation (4). Crucially, the approach never requires access to the posterior $p_{X|Y}(\cdot|\cdot)$. All we need is possibility of forward channel simulation, i.e., we assume that generation of samples $\bm{x}\sim p_{X}(\cdot)$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition are computationally feasible.
@@
  • Score-only MI estimation for nonlinear front-ends: We formalize a pipeline— DSM learning of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ Fisher integral representation—that applies to arbitrary deterministic $f$ as soon as $p_{Y_{t}}$ can be easily sampled.
  • Practical recipe: We detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral.
  • Validation and use-cases: We verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
In linear Gaussian channels, multivariate de Bruijn identity provides differential and integral representations of information quantities. For nonlinear $f$, however, posterior-free and normalization-free MI estimators remain limited. Building on DSM learning, our approach attaches directly to the marginal density $p_{Y_{t}}$, circumventing posterior computations and partition functions. Moreover, comparing with the conventional kernel density estimation (KDE) based estimators, our approach is robust to the high-dimensionality of the input space. The estimate of MI can be used for many applications such as design of coding systems, optimization of preprocessing, sensor design, and system optimization, and design of detection and decoding algorithms.
@@ II-A Notation
$J(U)\equiv\mathbb{E}_{\bm{u}\sim p_{U}(\cdot)}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right]$ (5)
@@ II-B Related works
For channels with memory, simulation-based methods have been developed to evaluate information rates without closed-form expressions. Arnold et al. [3] proposed a simulation-based framework that computes upper and lower bounds on the information rate by combining auxiliary finite-state channel models with forward–backward recursions and importance sampling; the method is broadly applicable to intersymbol interference (ISI) and finite-state channels where exact evaluation is infeasible. Building on this line, Sadeghi et al. [4] optimized these bounds over the choice of auxiliary models, yielding systematically improved estimates and a practical tool for design and benchmarking of channels with memory.
Another line of work estimates the MI directly from its definition via simulation-based plug-in estimators: one approximates $p_{Y}$ (or $h(Y)$) using KDE [5] or employs $k$-nearest neighbors (kkNN) estimators [6]. In the additive Gaussian setting, recent theory shows that plug-in estimation of $h(X+Z)$ is in fact optimal [7].
@@ II-C AWGN channel and score function
where $t$ represents the noise variance. The matrix $\mathbf{I}_{n}$ is the $n\times n$ identity matrix. Let $p_{Y_{t}}$ denote the marginal density of $Y_{t}$. The score function of $p_{Y_{t}}$ is defined by
$s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}).$ (7)
Throughout, we assume that sampling from $p_{X}$ is computationally feasible by forward sampling or Langevin sampling using the score $\nabla_{\bm{x}}\log p_{X}(\bm{x})$ (e.g., Rossky et al. [8], Welling and Teh [9]).
@@
$J(Y_{t})\equiv\mathbb{E}_{\bm{y}\sim p_{Y_{t}}(\cdot)}\bigl[\|s_{Y_{t}}(\bm{y})\|^{2}\bigr],$ (8)
where $\|\cdot\|$ denotes the Euclidean norm. Under standard regularity conditions, the following Tweedie identity [10, 11, 12] holds for $t>0$:
@@ II-D I–MMSE relation
Let $h(Y_{t})$ denote the differential entropy of $Y_{t}$. Consider the AWGN channel setup defined by (6). Then the (multivariate) de Bruijn identity gives
@@
The de Bruijn identity was discussed in [15, 16], which is proved by using the heat partial differential equation.
@@ II-F Denoising score matching (DSM) method
where this equivalence was shown by Vincent [18]. This means that minimizing the DSM loss function is equivalent to minimizing the expectation of the squared distance between the learned score function and the true score function, which is often called the Fisher divergence. This means that the unique minimizer of the DSM loss function satisfies $s_{\bm{\theta}}(\bm{y})\simeq\nabla_{\bm{y}}\log p_{\sigma}(\bm{y})$. Equivalently, since $\bm{v}-\bm{w}=\sigma\bm{\epsilon}$, one may implement the loss function as $(\bm{v}-\bm{w})/\sigma^{2}$:
@@
In practice, we use a neural network to parameterize the score model $s_{\bm{\theta}}(\cdot)$. Such a score model is also called a score network. A mini-batch training with stochastic gradient descent or other variants such as Adam is typically used to minimize the DSM loss function.
@@ III Estimating Mutual Information in AWGN Channels
In this section, we present a method to estimate the MI for AWGN channels described by
@@ III-A MI–Fisher relation
A sketch of proof is the following. Using the definition of MI,
$I(X;Y_{t})=h(Y_{t})-h(Z_{t})$ (21)
with $h(Z_{t})=(n/2)\log(2\pi e\,t)$, we have
@@
by de Bruijn identity. ∎
It can be remarked that, from the proposition, we immediately obtain the MMSE–Fisher relation
$\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ (23)
by the I–MMSE relation.
Note that Reeves et al. [21] investigated related identities in the context of linear Gaussian channels, highlighting its implications for non-Gaussian input distributions. Their work extends the I–MMSE framework beyond the classical Gaussian setting.
III-B Fisher Integral representations of mutual information
@@
$I(X;Y_{T})=$ $\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt.$ (24)
@@
The integral representation directly follows by combining the de Bruijn identity [15] with the I–MMSE relation [13, 14] as well.
@@ III-C Estimation of Fisher Information
Using DSM learning, we can estimate the score function of $Y_{t}$. With $p\leftarrow p_{X}$ and $\sigma^{2}\leftarrow t$, the learned score model can approximate the score of $Y_{t}=X+Z_{t}$:
$s_{\bm{\theta}}(\bm{y})\approx\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=s_{Y_{t}}(\bm{y}),$ (26)
where $s_{\bm{\theta}}(\cdot)$ can be implemented with a neural network with parameter vector $\bm{\theta}$. This enables estimation of $J(Y_{t})=\mathbb{E}_{\bm{y}\sim p_{Y_{t}}(\cdot)}\|s_{Y_{t}}(\bm{y})\|^{2}$ from samples of $\bm{x}\sim p_{X}$ and Gaussian noises.
@@
$=\mathbb{E}_{\bm{y}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{y})-s_{Y_{t}}(\bm{y})\,\right\|^{2}\right]+\text{const},$ (28)
where the constant term does not depend on $\bm{\theta}$. Hence, with a sufficiently rich function class for $s_{\bm{\theta}}$ (e.g., a universal neural network) and sufficiently many samples, any sequence of empirical minimizers achieves $\mathbb{E}_{\bm{y}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{y})-s_{Y_{t}}(\bm{y})\,\right\|^{2}\right]\rightarrow 0,$ i.e., the learned score converges to the true score in the Fisher divergence sense.
@@ III-D Gaussian input case for validation
In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$ where $P>0$ represents the signal power. It is well known [2] that $I(X;Y_{t})=(n/2)\log(1+P/t)$; see also [13] for the relation to MMSE.
@@
and $J(Y_{t})=\mathbb{E}\|s_{Y_{t}}(Y_{t})\|^{2}=n/(P+t)$. The remaining identities follow by substituting $J(Y_{t})$ into $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ and by $h(X|Y_{t})=h(Y_{t})-I(X;Y_{t})$ with $h(Y_{t})=(n/2)\log\big((2\pi e)(P+t)\big).$
@@ III-E Linear Gaussian channel case
IV Estimation of Mutual Information
Using the score-based estimation of the Fisher information, we can evaluate $I(X;Y_{T})$ by numerical integration. In this section, we present details of the method to estimate the MI for the AWGN channel family described by (6).
IV-A Score-based Estimation Method for MI
@@
Algorithm 1 Score-to-Fisher bridge: Estimation of MI
1: Prepare a logarithmically spaced grid over $t$ (details of the log grid are described in Subsection IV-B).
@@
3: DSM learning for $p_{Y_{t}}$: train the score model parameters by minimizing the empirical DSM loss:
@@
Any stochastic optimizer, such as Adam, can be applied to minimize the empirical loss function.
@@
$\widehat{J}_{t}=\frac{1}{N}\sum_{i=1}^{N}\big\|s_{\bm{\theta}}(\bm{y}_{i})\big\|^{2}$ (39)
@@
5: Estimating the MI via the Fisher integral representation:
@@
    where $\mathcal{T}$ is a training distribution over $t$ (e.g., logarithmically spaced as described later) and $\lambda(t)$ is a weight function, e.g., $\lambda(t)=t$.
In the case of Scheme B, the score network has two inputs: the channel output $\bm{y}$ and the noise level $t$.
IV-B Log-domain trapezoid integration
The integrand in the Fisher integral representation (4) of the MI is smooth and may have a long tail. The choice of the integration grid is critical for achieving high numerical accuracy. In this subsection, we present a geometric (log-spaced) grid for the integration. For $t\in[t_{\min},t_{\max}]$ with $M$ points, take $\log t$ to be equally spaced: